1.evaluate |2x+3|-|x-4|given that x<-2
2.solve the inequality (x^2-1)/(x^2-4)>=0
3.solve the inequality 1/(x-2)-1/(x-1)>=0

Answer by lwsshak3(6522)

1.evaluate |2x+3|-|x-4|given that x<-2
2.solve the inequality (x^2-1)/(x^2-4)>=0
3.solve the inequality 1/(x-2)-1/(x-1)>=0
..
1. |2x+3|-|x-4|=-(2x+3)-(x-4)=-2x-3-x+4=-3x+1
..
2.(x^2-1)/(x^2-4)>=0
write in factored form,
(x+1)(x-1)/(x+2)(x-2)>=0
Critical points: -2, -1, 1, 2
Show critical points on a number line.
Using test points,determine whether x-values in the following intervals make the function positive or negative.
(-infinity,-2),(-2,-1),(-1,1),(1,2),(2,infinity)
Starting from the right side, numbers >2 will make the(2,infinity)interval positive. Moving to the left, signs of the intervals will switch every time we go thru a critical point. (This occurs only because critical points are of multiplicity 1 or odd) Starting from the right,(2,infinity)is positive,(1,2) is negative,,(-1,1)positive,(-2,-1)negative,and (-infinity,-2)positive.
Solution:(-infinity,-2] U [-1,1] U [2,infinity). Brackets (]) are used to show that the end points are included in satisfying the function.
..
3.1/(x-2)-1/(x-1)>=0
combine into a single fraction.
(x-1)-(x-2)/(x-1)(x-2)>=0
x-1-x+2/(x-1)(x-2)>=0
1/(x-1)(x-2)>=0
critical points are 1 and 2
Put these points on a number line.
Following the same procedure as above,
Solution: (-infinity,1] U [2, infinity)